OpenStudy (anonymous):

Determine f′(x) for f(x)=(ln(3x))^9.

OpenStudy (calculusfunctions):

Have you learned derivatives of logarithmic functions?

OpenStudy (anonymous):

yea dont we use chain rule?

OpenStudy (anonymous):

ln u = u'/u

OpenStudy (calculusfunctions):

If\[y =c [\ln f(x)]^{n}\]then\[\frac{ dy }{ dx }=cn[\ln f(x)]^{n -1}\frac{ f \prime(x) }{ f(x) }\]because\[\frac{ d(\ln f(x)) }{ dx }=\frac{ f \prime(x) }{ f(x) }\]Note that in the above c is a constant. Does this help?

OpenStudy (anonymous):

i dont get it :C

OpenStudy (anonymous):

soo 9(ln(3x))^8(3/3x)

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

ohhh ok ty.